Jacarda Company makes a composting bin that is subject to wide seasonal variations in demand. Unit product costs are computed on a quarterly basis by dividing each quarter's manufacturing costs (materials, labor, and overhead) by the quarter's production in units. The company's estimated costs, by quarter, for the coming year given below:
Quarter First Second Third Fourth Direct materials $240,000 $120,000 $ 60,000 $180,000 Direct labour 96,000 48,000 24,000 72,000 Manufacturing overhead 228,000 204,000 192,000 216,000 Total manufacturing costs $564,000 $372.000 $276,000 $468,000 Number of units to be produced 80,000 40,000 20,000 60,000 Estimated unit product cost S 7.05 9.30 S 13.80 S 7.80
Management finds the variation in unit product costs to be confusing and difficult to work with. It has been suggested that the problem lies with manufacturing overhead, since it is the largest element of cost. Accordingly, you have been asked to find a more appropriate way of assigning manufacturing overhead cost to units of product. After some analysis, you have determined that the company's overhead costs are mostly fixed and therefore show little sensitivity to changes in the level of production.
Required:
1. The company uses a job-order costing system. How would you recommend that manufacturing overhead cost be assigned to production? Be specific, and show computations.
2. Recompute the company's unit product costs in accordance with your recommendations in (1) above.
Answer #1

1)

Company can use a predetermined overhead rate on the basis direct labour dollars. So that company can produce its units for a uniform price.

Predetermined Rate = Total Manufacturing Overhead Cost/Total Direct labor cost

Total Manufacturing Overhead = $228,000 +

$204,000 +

$192,000 +

$216,000

= $840,000

Total Direct Labor Cost = $96,000 +

$48,000 +

$24,000 +

$72,000

= $240,000

Predetermined Rate = $840,000/$240,000

= 3.5

Overhead Cost for Q1 :-

$96,000*3.5 = $336,000

Overhead Cost for Q2 :-

$48,000*3.5 = $168,000

Overhead Cost for Q3 :-

$24,000*3.5 = $84,000

Overhead Cost for Q4 :-

$72,000*3.5 = $252,000

2)

Q1 Q2 Q3 Q4
Direct Materials $240,000 $120,000 $ 60,000 $180,000
Direct Labor $ 96,000 $ 48,000 $ 24,000 $ 72,000
Manufacturing Overhead $336,000 $168,000 $ 84,000 $252,000
Total Manufacturing Cost $672,000 $336,000 $168,000 $504,000
Number of units Produced 80,000 40,000 20,000 60,000
Cost per unit $8.4 $8.4 $8.4 $8.4